OpenStudy (anonymous):

Given: ΔABC is a right triangle. Prove: a2 + b2 = c2 Right triangle BCA with sides of length a, b, and c. Perpendicular CD forms right triangles BDC and CDA. CD measures h units, BD measures y units, DA measures x units. The two-column proof with missing justifications proves the Pythagorean Theorem using similar triangles: Which is not a justification for the proof? Substitution Addition Property of Equality Transitive Property of Equality Distributive Property of Equality

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please, we have to apply the similitude criterions

OpenStudy (anonymous):

Statement Justification Draw an altitude from point C to Line segment AB Let segment BC = a segment CA = b segment AB = c segment CD = h segment DB = x segment AD = y y + x = c c over a equals a over y and c over b equals b over x a2 = cy; b2 = cx a2 + b2 = cy + b2 a2 + b2 = cy + cx a2 + b2 = c(y + x) a2 + b2 = c(c) a2 + b2 = c2

OpenStudy (michele_laino):

ok! I got since triangle BDC and triangle ADC are similar to triangle ABC, we can write: \[\frac{ a }{ y }=\frac{ c }{ a },\] and \[\frac{ b }{ x }=\frac{ c }{ b }\] from those equation, we can write: \[a ^{2}=cy, b ^{2}=cx\] so: \[a ^{2}+b ^{2}=cx+cy=c(x+y)=c ^{2}\]

OpenStudy (anonymous):

nvm I got it right it is transitive property of equality. thanks though

OpenStudy (michele_laino):

so I've used the first option, the second option, but not the third option

OpenStudy (michele_laino):

sorry I think that transitive property is used, because if a^2+b^2=c(x+y), and x+y=c, I can write: a^2+b^2=c^2

OpenStudy (anonymous):

it was transitive property I got it right. thanks for your help an explaining I really appreciate it!!(:

OpenStudy (michele_laino):

thank you!